OpenStudy (anonymous):

Help please!! I have no idea how to do this. Find the Percent increase that the following equation models. y=34(1.4)^x

jimthompson5910 (jim_thompson5910):

hint: rewrite 1.4 into the form 1+r

OpenStudy (anonymous):

so like y=34(1+r)^x ? Im horrible at math....

jimthompson5910 (jim_thompson5910):

if 1.4 = 1+r then r = ???

OpenStudy (anonymous):

i have no idea.. 1.4?

jimthompson5910 (jim_thompson5910):

1 + _____ = 1.4 what goes in the blank

OpenStudy (anonymous):

.4?

jimthompson5910 (jim_thompson5910):

good, 1 + 0.4 = 1.4

jimthompson5910 (jim_thompson5910):

that means 1.4 is in the form 1+r where r = 0.4

jimthompson5910 (jim_thompson5910):

that r = 0.4 means we have a 40% increase each time we increase x (x is a whole number)

OpenStudy (anonymous):

so thats the answer then?

OpenStudy (anonymous):

can you help me with one more question?

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (anonymous):

Its the other question i tagged you in